In the past week we made some progress with the Constant Force Drive (CFD). We found that one of the units of the Current to Voltage board had burned out. We also found that all of our spares have had a similar problem. We ordered an industrial drop-in part which does not quite fit the form factor for the electronics rack but did allow us to continue commissioning. We still have not determined why the boards were burned out but we will try to refurbish the parts and replace them as time allows. We also have one of our Telescope Operators now running the tracker. On a less positive note we found some metal fillings in the lower X encoder. This may simply be the race finding a new home under the weight of the tracker but we need some additional testing to determine if further damage may be occurring.
The mirror team was able to get all of the edge sensors back on the mirrors but we have run into significant communication problems with the CCAS alignment tower that did not allow us to fine align any of the mirrors in the array.
This week will be taken up with getting the communication problems with the CCAS alignment tower sorted out, investigating the problems with the lower X encoder, and trouble shooting the Current to Voltage boards.
We will also have contractors from Veliz Construction on site to start our big Remote Thermal Area project which actually includes a number of small projects including: Ducting heat from the Krabbenhut away from the telescope, ducting away and diluting acid vapor from the strip and wash room, installing the remote air conditioning and glycol chilling units, and installing the huge liquid nitrogen tank for the VIRUS instrument.